OpenStudy (anonymous):

An investment is worth $3518 in 1995. By 2000 it has grown to $5553. Let y be the value of the please help!!! investment in the year x, where x = 0 represents 1995. Write a linear equation that models the value of the investment in the year x.

OpenStudy (dape):

There are a few different ways to solve this, which class are you in?

OpenStudy (anonymous):

pre calc @dape

OpenStudy (anonymous):

y = 407x + 3518 i got this?

OpenStudy (dape):

Okay, so think like this, if I have say \(d\) amount of dollars the starting year, and the annual growth rate is let's say \(5\%\), then the next year I will have \[y(1)=1.05d\] The year after this, this whole lump of money will have grown by yet 5% again, so we have \[y(2)=1.05*y(1)=1.05*(1.05d)=1.05^2d\] Now, you probably see the pattern, after another year, we will multiply by 1.05 again to get a cube instead of a square and so on, so the general formula is: \[y(x)=d*r^x\] Where \(r\) is short for the growth rate and \(x\) is the number of years (2=squared, 3=cubed etc) and \(d\) is the starting money (this is the same as \(y(0)\)).

OpenStudy (dape):

Hmm, your answer could be right, depends on what they mean by linear equation. It could mean the more general property of linearity. In this case probably not.

OpenStudy (dape):

I think that yours is the correct answer, even if money does not grow in that way in the real world.

OpenStudy (anonymous):

@DAPE THANKS I got it right, i was confuse but thanks!